# GeoTikzBridge: Advancing Multimodal Code Generation for Geometric Perception and Reasoning

### 24 Mar 2026 v1

@@ 3.2 Iterative Self-Refinement for Image-to-Tikz Dataset Construction and Model Training
where $k=1,2,\ldots,K$, and $K$ denotes the maximum iteration number, which is set to $4$ in our experiments for a balance between cost and performance. The overall process for image-to-tikz data construction and model training consists of the following three steps:
@@
Sample Generation. Guided by the self-refinement paradigm, we select reliable samples from the constructed candidate dataset $\mathcal{D}_{\text{cand}}$ as potential additions to the training set by applying a CLIP score-based filtering strategy. The process begins with the generation of tikz pseudo-label set $\{\hat{C}\}$ for the geometric images, which are predicted by the model $M_{k}$. Then, the CLIP score (a normalized cosine similarity metric) $s$ between the image $\hat{I}$ rendered from the predicted tikz code $\hat{C}$ and the original image $I$ is computed. If $s$ exceeds a threshold $\tau$, the code $\hat{C}$ and the rendered image $\hat{I}$ would be considered as a reliable sample and incorporated in the self-refined set $\mathcal{D}_{k}^{R}$. $\mathcal{D}_{k}^{R}$ is defined as:
@@
The transformation strategy consists of two components: code transformation applied during the iterative dataset expansion, and image transformation applied during the model optimization (see Supplementary Material for details). Code transformation is operated on each tikz code in the self-refined set $\mathcal{D}_{k}^{R}$ at the $k$-th iteration, which randomly removes $1$ to $n$ lines (where $n$ does not exceed 40% of the total tikz code lines splitted with the semicolon). Each compilable modified code $\tilde{C}$ along with its rendered image $\tilde{I}$ serve as a sample pair in the transformation-based incremental dataset $\mathcal{D}_{k}^{T}$ at the $k$-th iteration. Denote the transformation operation as $\mathcal{T}(\cdot)$, $\mathcal{D}_{k}^{T}$ can be defined as:
@@
For the $k$-th iteration, the model $M_{k}$ is initialized from the previous model $M_{k-1}$ and optimized on the $k$-th training dataset $\mathcal{D}_{k}$. The training follows the causal auto-regressive modeling paradigm, where each code token in the tikz code sequence $c_{i}\in C$ is predicted conditioned on the image $I$ and all previous tokens $c_{<i}$. The corresponding training objective is defined as:
@@ 3.4 Training-Free Visual Reasoning in Geometric Problem-Solving
Ensure: Solution $A$ to the problem
@@ 4.1 Settings
Benchmarks and Metrics. We evaluate our models on five open-sourced benchmarks (DaTikZ[5], MathVista-GPS[26], GAOKAO-MM-Math[58], R-Bench[17], RBench-V[16]) and two in-house benchmarks (EDUBenchmark and GeoTikz-Instruct), focusing on mathematical visual reasoning. Following[36, 50], we adopt CLIP Score[33] and FID[33] for image-to-tikz task, along with fine-grained scores including MSE, SSIM, and PSNR for image-to-auxiliary-line task. Besides, we introduce a stricter metric, Compilation Success Rate, which is defined as the percentage of codes that can be successfully compiled and produce valid non-blank images, since most existing methods neglect the compilation errors that may occur in the code-to-image process. Details are introduced in Supplementary Material.
@@ F.1 Benchmark Details
five open-sourced benchmarks (i.e., DaTikZ, MathVista-GPS, GAOKAO-MM-Math, RBench, and RBench-V) and two in-house constructed benchmarks named EDUBenchmark and the constructed GeoTikz-Instruct are used for evaluation, most of which are focused on geometric mathematical reasoning.
@@
  • GAOKAO-MM-Math is a subset of GAOKAO-MM with 80 questions that focus on mathematical reasoning and is a Gaokao-based benchmark. The GeoTikzBridge-Base results are reported on its 142 images.

### 27 Mar 2026 v2

@@ 3.2 Iterative Self-Refinement for Image-to-Tikz Dataset Construction and Model Training
where $k=1,2,\ldots,K$, and $K$ denotes the maximum iteration number, which is set to 44 in our experiments for a balance between cost and performance. The overall process for image-to-tikz data construction and model training consists of the following three steps:
@@
Sample Generation. Guided by the self-refinement paradigm, we select reliable samples from the constructed candidate dataset $\mathcal{D}_{\text{cand}}$ as potential additions to the training set by applying a CLIP score-based filtering strategy. The process begins with the generation of tikz pseudo-label set $\{\hat{C}\}$ for the geometric images, which are predicted by the model $M_{k}$. Then, the CLIP score (a normalized cosine similarity metric) $s$ between the image $\hat{I}$ rendered from the predicted tikz code $\hat{C}$ and the original image II is computed. If $s$ exceeds a threshold $\tau$, the code $\hat{C}$ and the rendered image $\hat{I}$ would be considered as a reliable sample and incorporated in the self-refined set $\mathcal{D}_{k}^{R}$. $\mathcal{D}_{k}^{R}$ is defined as:
@@
The transformation strategy consists of two components: code transformation applied during the iterative dataset expansion, and image transformation applied during the model optimization (see Supplementary Material for details). Code transformation is operated on each tikz code in the self-refined set $\mathcal{D}_{k}^{R}$ at the $k$-th iteration, which randomly removes 11 to $n$ lines (where $n$ does not exceed 40% of the total tikz code lines splitted with the semicolon). Each compilable modified code $\tilde{C}$ along with its rendered image $\tilde{I}$ serve as a sample pair in the transformation-based incremental dataset $\mathcal{D}_{k}^{T}$ at the $k$-th iteration. Denote the transformation operation as $\mathcal{T}(\cdot)$, $\mathcal{D}_{k}^{T}$ can be defined as:
@@
For the $k$-th iteration, the model $M_{k}$ is initialized from the previous model $M_{k-1}$ and optimized on the $k$-th training dataset $\mathcal{D}_{k}$. The training follows the causal auto-regressive modeling paradigm, where each code token in the tikz code sequence $c_{i}\in C$ is predicted conditioned on the image II and all previous tokens $c_{<i}$. The corresponding training objective is defined as:
@@ 3.4 Training-Free Visual Reasoning in Geometric Problem-Solving
Ensure: Solution AA to the problem
@@ 4.1 Settings
Benchmarks and Metrics. We evaluate our models on five open-sourced benchmarks (DaTikZ[5], MathVista-GPS[26], GAOKAO-$M$-Math[58], R-Bench[17], RBench-V[16]) and two in-house benchmarks (EDUBenchmark and GeoTikz-Instruct), focusing on mathematical visual reasoning. Following[36, 50], we adopt CLIP Score[33] and FID[33] for image-to-tikz task, along with fine-grained scores including MSE, SSIM, and PSNR for image-to-auxiliary-line task. Besides, we introduce a stricter metric, Compilation Success Rate, which is defined as the percentage of codes that can be successfully compiled and produce valid non-blank images, since most existing methods neglect the compilation errors that may occur in the code-to-image process. Details are introduced in Supplementary Material.
@@ F.1 Benchmark Details
five open-sourced benchmarks (i.e., DaTikZ, MathVista-GPS, GAOKAO-$M$-Math, RBench, and RBench-V) and two in-house constructed benchmarks named EDUBenchmark and the constructed GeoTikz-Instruct are used for evaluation, most of which are focused on geometric mathematical reasoning.
@@
  • GAOKAO-$M$-Math is a subset of GAOKAO-$M$ with 80 questions that focus on mathematical reasoning and is a Gaokao-based benchmark. The GeoTikzBridge-Base results are reported on its 142 images.